OpenStudy (anonymous):

What number must be added to the polynomial to complete the square? x2+13x

OpenStudy (anonymous):

here we go... half of 13 is \[\frac{13}{2}\] \[(\frac{13}{2})^2=\frac{169}{4}\] and that is your answer

OpenStudy (anonymous):

not that, it is just that by now it should be routine. half of "x" term is ____ then _____ squared is "answer"

OpenStudy (anonymous):

so you can do it pretty much in your head, as long as you know a couple squares

OpenStudy (anonymous):

complete the square for \[x^2+bx\] answer is \[x^2+bx+(\frac{b}{2})^2\]

OpenStudy (anonymous):

capice?

OpenStudy (anonymous):

but what about this one,,,,x^2+x